‘n’ litres of oil was poured into a tank and it was still e% empty. How much oil must be poured into the tank in order to fill it to the brim.

Answer by stanbon(57387)

‘n’ litres of oil was poured into a tank and it was still e% empty. How much oil must be poured into the tank in order to fill it to the brim.
------------------------------------
The n litres occupies (100-e)% of the tank.
--------------------------
Proportion:
Let x litres be the capacity of the tank
n/(100-e)% = x/100%
--------
x = n/(100-e)%
------------
The tank already has n litres in it.
You need [n/(100-e)%]-n = [100n/(100-e)]-n
= [100n-n(100-e)]/(100-e)
= ne/(100-e) litres to fill the tank
==================
Cheers,
Stan H.
================